5z-2y=-1
Y intercept is -1 x is the rise is 2 and the run is 1 so go up 2 and ove 1 and you will reach the x axis. do same for line 2


Found 2 solutions by richwmiller, miller95:
Answer by richwmiller(17219)

multiply the first by 2
10x-4y=-2
x+4y=35
add
11x=33
x=3
3+4y=35
4y=32
y=8
(3,8)
To plot the lines do, one line at a time.
let y=0
x+0=35
x=35
(35,0)
slope is -1/4
y goes down 1 and x goes up 4
39,-1
5x-2y=-1
letx=0
-2y=-1
y=1/2
let y=0
5x=-1
x=-1/5

Answer by miller95(1)

5z-2y=-1
Y intercept is -1 x is the rise is 2 and the run is 1 so go up 2 and ove 1 and you will reach the x axis. do same for line 2